# Graph Reinforcement Learning in Power Grids: A Survey

### v1

@@ 6 Other Applications
Model RL Action GNN Focus
@@
Communication
Islam et al., (2023) Q-Learning Routing, set queue service rate spectral GCN reduce end-to-end latency, packet loss rate
@@ 6.1 Energy Market
GRL opens new possibilities in the energy market, especially in decentralized bidding or direct trading among entities. Traditionally, bidding strategies are centrally managed, requiring full information on all generation units. This is often unavailable due to privacy concerns and results in large-scale, computationally expensive problems. Distributed decision-making, using multi-agent RL and GNNs has the potential to provide efficient and scalable solutions.
@@
Rokhforoz et al., (2023) consider a two-level optimization: first, each unit sets a bidding price; second, the market operator determines the market price. The goal is to maximize each unit’s profit. They propose a multi-agent actor-critic algorithm, where each unit’s actor network selects a bidding strategy based on previous market data. After the market price is set, rewards are calculated. The critic network, a GNN, updates the bidding strategy based on the grid topology. Experiments on the IEEE-30-bus and 39-bus systems show that the GNN approach outperforms the baseline using an MLP-based critic, particularly with varying generation capacities. The GNN also demonstrated better transfer capability when tested across different systems.
In the approach by Lee et al., 2022a , energy is traded directly between prosumers without an intermediary market. The setup includes multiple nanogrids, an information network, and a business network for trading. The proposed RL algorithm learns trading strategies to minimize maximum load and maximize renewable integration, including power from discharging EVs. The agent’s actor and critic are hybrid models combining a GCN with a Bi-LSTM to process time series data on prosumer consumption and production. The model inputs include cluster demand, renewable supply, system price, and demand response. Actions are buy, sell, or hold, and rewards are based on a rule-based baseline. Multi-objective optimization includes load shifting. The authors compare various RL methods, including DQN, Bi-LSTM, and PPO, using a nanogrid with real usage data. The PPO GCN-Bi-LSTM approach achieves the lowest electricity cost and performs better than other methods, significantly reducing average electricity costs with P2P trading.
@@ 6.2 Power Communication Networks
Islam et al., (2023) develop a prioritized routing strategy considering grid states. They address two packet types: periodic packets with fixed schedules and emergency packets needing low latency. The goal is to reduce end-to-end latency, packet loss, and unmet latency requirements using software-defined networks that adapt dynamically to grid conditions.
Two Q-Learning RL algorithms are trained: one for routing paths and another for queue service rates to minimize congestion. The first agent selects feasible paths, while the second predicts queue service rates at switches. Rewards are based on the difference between switch capacity and queue state in order to accelerate queue emptying. A GNN predicts future grid states to inform the queue service rate agent, though the GNN is trained separately from the RL agent. The model, using spectral GCN with Chebyshev polynomials, is trained on IEEE grid traffic data. Experiments on the cyber layers of the IEEE-14 and 39-bus systems show the approach’s effectiveness in managing grid communication through message exchanges between devices and control centers.
@@ 6.3 Electric Vehicles Applications
The rapid growth of electromobility has posed challenges with increasing demands for electrical energy and straining electricity networks, particularly with variable load behaviors and decentralized renewable energy supplies. Efficient infrastructure use has become critical. DRL has found applications across several domains including proof-of-concepts for charging management Bayani et al., (2022); Sadeghianpourhamami et al., (2020); Li et al., (2020); Silva et al., (2020), station recommendation Xing et al., 2023b ; Xu et al., 2022c , navigation Xing et al., 2023c ; Xu et al., 2022c ; Xing et al., 2023b , and pricing optimization Zhang et al., (2022). These applications optimize energy distribution, route EVs efficiently, and dynamically adjust pricing to balance operator revenue with infrastructure efficiency.
We focus on GRL approaches and highlight two studies that include power grid constraints in optimizing charging processes, while noting the broad applications of GRL in various EV charging scenarios.
Xu et al., 2022c address the demands on fast charging stations caused by the increasing number of EVs, which affect coupled power-transportation networks. They propose a multi-objective system-level method to dynamically allocate EVs to suitable stations using DRL. This method considers the interests of EV owners, charging stations, traffic networks, and power grids, ensuring real-time responsiveness and efficient allocation of fast charging resources. EVs send charging requests when their state of charge drops below a threshold, and an agent recommends a charging station and guides them using Dijkstra’s algorithm. The multi-objective function optimizes EV wait times, charging station service balance, traffic congestion, and power grid voltage deviation.
Similarly, Xing et al., 2023b introduce a Bi-Level GRL approach for charging and routing in Transportation Electrification Coupled Systems. They use a bi-level problem formulation, with both levels employing a GNN-based feature extraction layer and a Rainbow-architecture DRL block for decision-making. The upper level recommends a charging station, while the lower level handles EV travel and route selection. Unlike Xu et al., 2022c , this approach uses a DRL agent for route planning instead of a conventional algorithm.
@@
GNNs leverage the inherent graph structures of transportation systems and power grids. Xu et al., 2022c designs a graph structure that captures connections between charging stations, roads, and power nodes. Charging stations bridge the traffic network and power grid, connecting to traffic nodes (midpoints of road sections) and power nodes based on geographical and power supply relations. The adjacency matrix of the charging network represents connections between traffic nodes (TN), power nodes (PG), charging stations (CS), and EV nodes. A unified expression method with type-specific transformation matrices projects features into a shared space, and GATs extract meaningful features from neighboring nodes using attention mechanisms. These learned representations are integrated with EV features for input to a DRL agent.
Similarly, Xing et al., 2023b utilizies GATs and introduces an instantaneous adjacency matrix for connections among EVs, CSs, TNs, and PGs, with smaller matrices representing different relationships (e.g., EV-TN, CS-TN). A node feature matrix stores energy and information features, with different rows representing system elements.
RL algorithms
When recommending a fast charging station, the decision is quick, but actual completion occurs after the EV finishes charging, causing significant delays. This delay, coupled with system complexity and unpredictability, makes forecasting the entire charging process challenging, impacting standard RL methods. To address this, Xu et al., 2022c introduce the double-prioritized DQN($\lambda$) method, integrating $\lambda$-return and experience replay to handle action delay and uncertainty. They use a small cache in replay memory for efficient training, prioritize crucial transitions, and employ an attention-based method to construct a cache that maintains high-quality training transitions. Additionally, a strategy regulates boundary actions to prevent detrimental decisions.
In contrast, Xing et al., 2023b employ a bi-level approach to address the credit assignment problem. The upper level selects charging stations less frequently, focusing on choosing a target CS, while the lower level deals with path navigation by selecting road segments. Rewards in each level consider charging costs, battery loss, time allocation, energy consumption, travel time, and penalties for exceeding voltage limits. The upper level interacts with charging stations and power grids, while the lower level interacts with traffic nodes after determining the target station.
@@
Xu et al., 2022c use a joint power-transportation simulation platform with an IEEE 33-node distribution network and a 25-intersection traffic network to validate their method. They optimize traffic conditions, user experience, and power grid stability, demonstrating superiority over distance-based methods across scenarios, even under charging station queue limitations. Training a user-oriented graph DQN($\lambda$) agent shows significant long-term benefits and user experience improvement. Combining GATs and DQN($\lambda$) outperforms other models in training, confirming its effectiveness in complex decision-making. However, a comparison of an MLP-DQN($\lambda$) remains to be performed in order to determine the effect of the GAT representation without any effects from the long term $\lambda$-return.
Xing et al., 2023b evaluate their Bi-Level GRL approach for Electric Vehicle Fleet Charging Guidance using real-size transportation-electrification data. They achieve 10.08% cost reduction and 16.45% time savings for owners compared to realistic scenarios without guidance. Compared to traditional methods, their approach lowers average total cost by 8.96% (SDP) and 4.73% (DRL), highlighting its efficiency in reducing charging costs and travel time. They suggest learning the GNNs weights directly with RL for enhanced robustness and scalability.
@@ Discussion
Xu et al. Xu et al., 2022c demonstrate GRL’s ability in dynamically allocating resources for fast charging stations, emphasizing real-time responsiveness and multi-stakeholder considerations. Their approach highlights sequential decision-making’s role in balancing objectives across transportation and power networks. In contrast, Xing et al. Xing et al., 2023b present a Bi-Level GRL approach, focusing on efficient charging and routing coordination through GNNs. Their method excels in reducing charging costs and travel time, demonstrating GRL’s potential in real-world scenarios. While both employ GAT architecture, further comparisons of GNN architectures and robustness to noisy data are essential for future research, in order to further GRL’s applicability in sustainable transportation and energy infrastructure optimization.
@@ 7 Conclusion and Outlook
The exploration of DRL and GRL approaches in addressing the challenges of power grid control, voltage management, and other related use cases such as EV charging management reveals significant advancements and promising avenues for future research and applications. Across the transmission grids, distribution grids, and EV charging domains, these approaches demonstrate innovative solutions that leverage machine learning techniques to optimize system performance, enhance resilience, and facilitate the integration of renewable energy sources and electric transportation.
@@
The pivotal role GNNs in RL approaches cannot be overstated. GNNs serve as the backbone for feature extraction from complex network structures, enabling the effective representation of power grid topologies. Their ability to capture spatial dependencies and relational information among grid components facilitates informed decision-making in multiple grid control applications. Moreover, the versatility of GNN architectures allows for customization to specific application requirements, ranging from spectral and spatial GCNs, to GATs, to GraphSAGE and beyond. Despite their widespread adoption in the presented papers, further research is needed to systematically evaluate and compare the suitability of GNN architectures for various control tasks within the power system domain. GNNs have shown to achieve better results than fully connected neural networks on unseen grid topologies and differing networks sizes.
@@
In real-world power systems, data quality issues such as noisy sensor measurements and missing data are common due to factors like sensor malfunctions, communication failures, and data transmission errors. These data imperfections can significantly impact the performance and reliability of RL models. GNNs excel at capturing complex spatial dependencies and relational information from graph-structured data, allowing them to effectively distinguish meaningful patterns even in the presence of noisy data. By leveraging the inherent connectivity and topology of the underlying network, GNNs can smooth out noise and extract robust representations that encode relevant features for decision-making tasks. While a few of the presented approaches validated their GRL methods against noisy measurements, more research is needed to establish robustness across diverse datasets and real-world conditions.
@@
Although different GNN architecture have been used across the approaches, the models are mostly very shallow with only 2-3 layers. As pointed out in Ringsquandl et al., 2021b , the graphs structure of GNNs differs from standard benchmark datasets, which makes well-established models not fully suitable to learning tasks on power grids. In fact, they find out that much deeper GNNs perform better, since they can capture long-range dependencies in graphs with a low cluster and high diameter as it it the case for power grids. Therefore, deeper architectures are an interesting topic of research especially in case of large-scale grids.
As for the RL methods, the research landscape for power grid control reveals a lack of application of state-of-the-art RL innovations. On the one hand, despite few efforts utilizing MCTS, the application of model-based RL methods remains limited, further exemplified by the absence of GNN based approaches. On the other hand, contemporary model-free RL innovations such as BBF Schwarzer et al., (2023) have yet to be explored in this context, which underscores a significant gap in current research.
@@
Another promising direction for future research is the incorporation of domain knowledge. In the discussed approaches, it had been used to identify relevant grid connections or components to reduce the actions space or augment the graph to introduce meaningful edge weights and enable faster information diffusion. Further, the customized graph shift operators proposed by several approaches have shown to improve the performance in GNNs. Making use of the known relations and coherence in power grids, therefore, should be investigated further. Additionally, further research is needed to explore the incorporation of physical properties, such as power flow constraints, into agent training. This could be achieved, for example, through the utilization of physics-informed GNNs.
@@
Variation in reward functions reflect the trade-offs involved in optimizing multiple objectives and highlights the ability of RL to handle complex, multi-dimensional optimization problems. Multi-objective RL offers flexibility by allowing the incorporation of various objectives into the reward function. However, balancing conflicting objectives remains a challenge, as optimizing one objective may have adverse effects on others. In future research, there is a pressing need to explore approaches where agents are capable of providing multiple suggestions instead of just a single solution. This could involve presenting a range of solutions along the Pareto front, showcasing the trade-offs between conflicting objectives, or offering alternatives that prioritize different objectives. By offering multiple suggestions, agents can enhance the trustworthiness and transparency of their recommendations, allowing users to gain a deeper understanding of the decision-making process. Moreover, presenting a variety of options empowers users to make more informed decisions that align with their preferences and operational constraints.
@@
In the transmission grid domain, many approaches focus on smaller grid sizes. There is potential to extend these techniques to larger grids. Investigating how these methods scale with grid size and complexity would be valuable for real-world deployment. Furthermore, the grid sizes evaluated in many of the presented papers are typically smaller than real-world grids, even for the larger test grids. Action space reduction, and exploration in constrained environments underscore the need for further methodological advancements, particularly as the power grid size increases. Handcrafted rules and rule-based algorithms have been successfully integrated into the different GRL solutions. The synergy between rule-based decisions and GRL methodologies showcases the importance of leveraging domain knowledge to achieve effective grid control.
@@
Notably, it’s important to point out the inherent bias towards the grid2op framework, which provides the data, training and testing environments for many approaches. While grid2op incorporates many real-world aspects, there’s a need to optimize graph representations for grid processing, which can significantly influence the performance of GNNs. Addressing the stochasticity in the grid2op environment is also crucial, as it significantly impacts agent performance. Averaging over multiple seeds is essential for reliability, yet often overlooked in some approaches. Transitioning GRL approaches from simulation environments like grid2op to real-world grid operation poses significant challenges, including computational efficiency, scalability, and integration with existing control systems.
For the distribution grid, the discussion surrounding voltage control and grid management GRL techniques underscores the diverse approaches and considerations within this domain. A central challenge lies in the design of reward functions, which typically prioritize voltage deviations while also accommodating factors like renewable energy integration and power loss. The choice of actuators, primarily focusing on PV inverters but also considering Energy Storage Systems and generator adjustments, highlights the need for adaptable control strategies. Multi-agent setups for zoned distribution grids present an interesting avenue, with varying approaches to incorporating global system knowledge. The implementation of GNNs varies widely, reflecting the ongoing exploration of their effectiveness in handling noisy or missing data. Despite consistent graph representations, there is significant diversity in GNN architectures. Evaluations span a range of grid sizes and metrics, demonstrating the scalability, stability, and real-time deployment capabilities of GRL-based frameworks. Additionally, further investigation into the augmentation of graph structures and the integration of temporal information could yield valuable insights for improving the performance and robustness of GRL-based approaches in distribution grid control.
The transition to fully autonomous systems must be approached cautiously, with a focus on developing decision support systems that augment human expertise rather than replacing it entirely. Therefore, exploring methods for enhancing the explainability of DRL approaches to ensure transparency and trustworthiness in the decision-making process remains crucial for the future. Besides explaining model decisions, assessing model uncertainty, and verifying model behavior under various conditions can help enhance transparency and confidence in DRL-based grid operation. The integration of GNNs into RL algorithms holds significant promise, since leveraging the inherent graph structure of power systems topologies can offer a more intuitive understanding of complex interactions. Architectures like GATs introduce interpretability through mechanisms such as attention coefficients, enabling the quantification of the importance of neighboring nodes. This does not only help to enhance model transparency but also facilitates more informed decision-making within RL frameworks.

### v3

@@ 6 Other Applications
This chapter explores GRL approaches for the related applications of new energy markets, communication networks for power grids, and EV charging scheduling. We consider only approaches that take into account the underlying power grid structure and constraints. Tab. 4 provides an overview of the RL method, action type, GNN architecture, grid size, and overall focus of the approaches analyzed.
@@
In the left column, Comm. stands for Communication Networks and in the column Action, CS refers to charging station.
Approach RL Action GNN Focus/ Unique Feature
@@
Comm. Islam et al., (2023) Q-Learning Routing, setting queue service rate Spectral GCN Reduce end-to-end latency, packet loss rate
@@ 6.1 Energy Market
GRL opens new possibilities in the energy market, especially in decentralized bidding or direct trading between entities. Traditionally, bidding strategies are centrally managed, requiring full information on all generation units. This is often infeasible due to privacy concerns and results in large-scale, computationally expensive problems. Distributed decision-making, using multi-agent RL and GNNs, has the potential to provide efficient and scalable solutions.
@@
Approaches
Rokhforoz et al., (2023) propose a two-level optimization as follows: first, each unit sets a bidding price; second, the market operator determines the market price. The goal is to maximize each unit’s profit.
The rewards are calculated based on the determined market prices. The critic network, a GNN, updates the bidding strategy based on the grid topology. Experiments on the IEEE 30-bus and 39-bus systems show that the GNN approach outperforms the baseline using an MLP-based critic, particularly under varying generation capacities. When tested across different systems, the GNN also demonstrated better transfer capability.
In the approach by Lee et al., 2022a , energy is traded directly between prosumers without an intermediary market. The setup includes multiple nanogrids, an information network, and a business network for trading. The proposed RL algorithm learns trading strategies to minimize maximum load and maximize renewable integration, including power from discharging EVs. The agent’s actor and critic are hybrid models combining a GCN with a Bi-LSTM to process time series data on prosumer consumption and production. The model inputs include cluster demand, renewable supply, system price, and demand response. The actions are either buy, sell, or hold. The reward is based on a rule-based baseline, and multi-objective optimization includes load shifting. The authors compare various RL methods, including DQN, Bi-LSTM, and PPO, using a nanogrid with real usage data. The PPO GCN-Bi-LSTM approach achieves the lowest electricity cost and performs better than other methods, significantly reducing average electricity costs with P2P trading.
@@ 6.2 Power Communication Networks
The study in Islam et al., (2023) addresses packet routing and presents a prioritized strategy considering grid states. They distinguish two types of packets: periodic packets with fixed schedules and emergency packets needing low latency. The goal is to reduce end-to-end latency, packet loss, and breach latency requirements using software-defined networks that adapt dynamically to grid conditions.
Two Q-Learning RL algorithms are trained: one for routing paths and another for queue service rates to minimize congestion. The first agent selects feasible paths, while the second predicts queue service rates at switches. Rewards are based on the difference between switch capacity and queue state in order to accelerate queue emptying.
A GNN predicts future grid states to inform the queue service rate agent, though the GNN is trained separately from the RL agent. The model, using spectral GCN with Chebyshev polynomials, is trained on IEEE grid traffic data. Experiments on the cyber layers of the IEEE-14 and 39-bus systems show the approach’s effectiveness in managing grid communication through message exchanges between devices and control centers.
@@ 6.3 Electric Vehicles Applications
The rapid growth of electromobility is challenging the grid infrastructure, as it increases electricity demand and introduces variable loads. In this context, DRL has been studied for charging management Bayani et al., (2022); Sadeghianpourhamami et al., (2020); Li et al., (2020); Silva et al., (2020), station recommendation Xing et al., 2023b ; Xu et al., 2022c , navigation Xing et al., 2023c ; Xu et al., 2022c ; Xing et al., 2023b , and pricing optimisation Zhang et al., (2022). These applications optimize the allocation of electricity, the pricing, and the routing of EVs. We concentrate on those GRL approaches that consider the underlying power grid and its constraints.
RL algorithms
The study in Xu et al., 2022c tackles the increasing demands of fast charging stations. They propose a multi-objective DRL method to dynamically allocate EVs to stations, considering EV owners, charging stations (CS), traffic nodes (TN), and power grid nodes (PG). The agent recommends CSs and guides EVs using Dijkstra’s algorithm, optimizing waiting times, service balance, traffic congestion, and grid voltage deviation. The recommendation of a CS is fast, but the full process is complete only once the EV finishes charging. The double-prioritized DQN($\lambda$) method is introduced to address this delay and unpredictability. It integrates $\lambda$-return and experience replay with a small buffer to improve efficiency. During training, high-quality samples are prioritized using an attention mechanism, along with a strategy to regulate boundary actions.
Similarly, Xing et al., 2023b present a Bi-Level GRL approach for charging and routing in Transportation Electrification Coupled Systems. Using a Rainbow-architecture DRL block, the upper level recommends CSs, while the lower level selects routes with a DRL agent. This bi-level approach addresses credit assignment by having the upper level select charging stations less frequently, focusing on the target CS, and the lower level handle path navigation. Rewards consider charging costs, battery loss, time allocation, energy consumption, travel time, and voltage limit penalties, with the upper level interacting with charging stations and power grids and the lower level with traffic nodes.
@@
GNNs leverage the inherent graph structure of transportation systems and power grids. Therefore, Xu et al., 2022c design a graph structure based on the physical properties. The CSs connect to TNs, and PGs are based on geographical and power supply relations. A unified expression method with type-specific transformation matrices projects features into a shared space, and GATs extract meaningful features. These learned representations are integrated with EV features for input to a DRL agent.
Similarly, Xing et al., 2023b utilize GATs and introduce an instantaneous adjacency matrix for connections among EVs, CSs, TNs, and PGs, with smaller matrices representing different relationships. Node features store energy and information features.
@@
The approach in Xu et al., 2022c is validated using a power-transportation simulation platform with an IEEE 33-node distribution network and a 25-intersection traffic network. They optimize traffic, user experience, and grid stability, outperforming distance-based methods even with charging station queue limits. Training a user-oriented graph DQN($\lambda$) agent shows long-term benefits and improved user experience. Combining GATs and DQN($\lambda$) improves training and decision-making, though a comparison with MLP-DQN($\lambda$) would be needed to investigate the impact of GATs.
On the other hand, Xing et al., 2023b evaluate their Bi-Level GRL approach using real transportation-electrification data, achieving a 10.08% cost reduction and 16.45% time savings for owners. Compared to other methods, their GRL approach lowers the average total cost by 8.96% (distance-based) and 4.73% (DRL), demonstrating its efficiency. They recommend learning GNNs weights directly with RL for better robustness and scalability.
@@ Discussion
The authors of Xu et al., 2022c demonstrate the effectiveness of GRL in dynamic resource allocation for charging stations, emphasizing real-time responsiveness and multi-stakeholder considerations. Their approach highlights the role of sequential decision-making in balancing objectives across transportation and power networks. In contrast, Xing et al., 2023b focuses on efficient charging and routing coordination through GNNs. Their method aims to reduce charging costs and travel time, demonstrating the potential of GRL in real-world scenarios. While both employ GAT architectures, other GNN architectures and robustness to noisy data should be investigated more closely to confirm the applicability of GRL in real transportation and energy infrastructures.
@@ 7 Conclusion and Outlook
The exploration of GRL approaches to the challenges of power grid control, voltage management, and other related use cases reveals significant advances and promising avenues for future research and applications. Across these applications, GRL approaches demonstrate solutions that leverage the inherent graph structure to optimize performance, enhance resiliency, and facilitate the integration of renewable energy sources.
@@
The role of GNNs in DRL approaches cannot be overstated. They enable effective representation of power grid topologies by extracting features from complex network structures and capturing spatial dependencies. This facilitates informed decision-making in grid control applications. Their versatility in terms of architecture allows them to be adapted to specific needs. Despite their use in the literature, further research is required to evaluate and compare architectures for different power system control tasks. GNNs show to be more effective than fully connected neural networks on unseen network topologies and differing grid sizes.
@@
In real power grids, data quality issues such as noisy sensor readings and missing data are common due to sensor malfunctions, communication failures, and transmission errors that impact RLs performance. GNNs excels at capturing complex spatial dependencies, enabling them to identify meaningful patterns even in noisy data. By aggregating neighbouring information they smooth noise and extract robust embeddings. Although some approaches have validated GRL in that context, more research is required to ensure robustness for real-world conditions.
@@
Although different GNN architectures have been used, most models are shallow with only 2-3 layers. As noted in Ringsquandl et al., 2021b , the graph structure of GNNs differs from standard benchmark datasets, making established models less suitable for power grids. They found that deeper GNNs perform better by capturing long-range dependencies in low-cluster, high-diameter graphs such as power grids. Thus, specialized architectures are a promising research topic, especially for large-scale grids. Additionally, optimizing graph representations for grids is necessary to improve GNN performance as it significantly influences the performance of a GNN.
On the other hand, RL methods for power grid control do not use state-of-the-art methods. Model-based RL methods, including MCTS, are scarcely used, with no GNN-based approaches combined with more sophisticated methods. In addition, contemporary model-free RL innovations such as BBF Schwarzer et al., (2023) remain unexplored, highlighting a significant research gap.
@@
Another promising direction for future research is the incorporation of domain knowledge. In the discussed approaches, relevant grid connections or components have been identified to reduce the action space, to augment the graph with meaningful edge weights or allow for faster information diffusion. Furthermore, customized graph shift operators proposed by several approaches have improved the performance in GNNs. Therefore, exploiting the known relationships in power grids should be further investigated. In addition, further research should incorporate physical properties, such as power flow constraints, into agent training, for example, through physics-informed GNNs.
@@
The variation in reward functions reflects the trade-offs in optimizing multiple objectives, demonstrating RL’s ability to handle complex, multidimensional problems. Multi-objective RL allows multiple objectives to be incorporated into the reward function. However, balancing conflicting objectives is challenging because optimizing one may negatively affect others. Future research should explore agents that provide multiple suggestions rather than a single solution. This could include presenting a range of solutions along the Pareto front, highlighting trade-offs, or offering alternatives that prioritize different goals. This would increase trustworthiness and transparency and allow users to make more informed decisions.
@@
For transmission grids, almost all approaches focus on grids that are smaller than real-world grids. Extending them to larger grids and investigating their scalability with grid size and complexity is crucial for real-world deployment. In this context, advances in action space reduction and exploration in constrained environments could help. Hand-crafted rules and rule-based algorithms have been successfully integrated into various GRL solutions, demonstrating the importance of leveraging domain knowledge for effective grid control.
@@
It is important to note the inherent bias of the grid2op framework, which provides data, graph representation, training, and testing environments for many approaches. While grid2op incorporates many real-world aspects, it is an abstraction from reality. Transitioning GRL approaches from simulation to real-world grid operations poses challenges, including computational efficiency, scalability, and integration with existing control systems. Addressing stochasticity in the grid2op environment by averaging over multiple seeds is also critical, as it significantly affects agent performance. However, it is often overlooked.
In the distribution grid, voltage control and grid management GRL techniques show different approaches. Designing reward functions is challenging because of the need to balance voltage deviations, renewable energy integration, and power loss. The variety of actuators, like PV inverters or ESS, highlights the need for adaptable control strategies. Multi-agent setups for zoned grids have shown promising results but are rarely investigated so far. Furthermore, it can be stated that the challenges in terms of scalability, real-time deployment, GNN architecture, and graph representations described above also apply to the distribution network.
The deployment of RL-based techniques and the transition to fully autonomous systems requires an emphasis on decision support systems that augment rather than replace human expertise. Enhancing the explainability of DRL approaches and assessing model uncertainty and behavior under varying conditions is, therefore, critical to ensuring transparency and trust in decision-making. Using GNN architectures such as GATs introduces interpretability by using attention coefficients to highlight important nodes. However, an analysis of the coefficients is missing in the presented approaches. The integration of GNNs into RL algorithms improves robustness to noise and missing data, as well as transferability. This integration contributes to model transparency and supports more informed decision-making.